OpenStudy (anonymous):

Help!! Solve the triangle,if possible.round to the nearest hundredth a=28ft b=31ft c=45ft

OpenStudy (mertsj):

Use the law of cosines.

OpenStudy (anonymous):

give me more hint please

OpenStudy (mertsj):

\[c^2=a^2+b^2-2ab \cos C\]

OpenStudy (anonymous):

i got lost :/

OpenStudy (anonymous):

39,060 ft.

OpenStudy (anonymous):

more hint

OpenStudy (mertsj):

\[45^2=28^2+31^2-2(28)(31)\cos C\] \[2025=784+961-2(28)(31)\cos C\] \[\frac{2025-784-961}{-1736}=\cos C\]

OpenStudy (anonymous):

Look up the law of cosines

OpenStudy (mertsj):

Find C Then repeat the process to find B

OpenStudy (anonymous):

Nevermind